Over a certain region of space, the electric potential is V = 2x - 5x2y + 2yz2. Find the expression for the x component of the electric field over this region. (Use the following as necessary: x, y, and z.)

Ex =  


Find the expression for the y component of the electric field over this region.

Ey =   

Find the expression for the z component of the electric field over this region.

Ez =  


What is the magnitude of the field at the point P, which has coordinates (2, 0, -4) m?
N/C
Answer #1

The relation between electric field (E) and electric potential (V) is given by

vec{E} = - , grad , (V) = - left ( hat{i} , rac{partial }{partial x} + hat{j} , rac{partial }{partial y} + hat{k} , rac{partial }{partial z} ight ) V

Given,

  V = 2x-5x^{2}y + 2yz^{2}

So,

vec{E} = - , grad , (V) = - left ( hat{i} , rac{partial }{partial x} + hat{j} , rac{partial }{partial y} + hat{k} , rac{partial }{partial z} ight ) (2x-5x^{2}y + 2yz^{2})

     = - left ( hat{i} , rac{partial (2x-5x^{2}y + 2yz^{2})}{partial x} + hat{j} , rac{partial (2x-5x^{2}y + 2yz^{2})}{partial y} + hat{k} , rac{partial (2x-5x^{2}y + 2yz^{2})}{partial z} ight )

  = - left ( hat{i} , (2 - 10 xy)+ hat{j} , (-5x^{2} + 2z^{2})+ hat{k} , (4yz) ight )

  = hat{i} , ( 10 xy - 2)+ hat{j} , (5x^{2} - 2z^{2})+ hat{k} , (-4yz)

We know that  

vec{E} = hat{i} , E_{x} + hat{j} , E_{y} + hat{k} , E_{z}

Comparing these two equations we get ,

E_{x} = 10 xy -2

  E_{y} = 5x^{2} - 2z^{2}

E_{z} = -4yz

At point P(2 , 0 , -4) m,

Er = 10 x 2 x 0-2=-2 N/C

E, 5(2)--2(-1)--20-32 =-12 N/C

E,--1 x 0 x (-4) = 0 N/C

So, magnitude of field at point P is

E = left | vec{E} ight | = sqrt{E_{x}^{2} + E_{y}^{2} + E_{z}^{2}}

  V/(-2)2 + (-12)2 + (O)2-V148 = 12.165 N/C

For any doubt please comment and please give an up vote. Thank you.